OpenStudy (anonymous):

find the numerical value of sin3pi/2 +pi/4

OpenStudy (anonymous):

Use trig identity: sin (a + b) = sin a*cos b* + sin b* cos a y = sin (3Pi/2 + Pi/4) = sin 3Pi/2 cos Pi/4 + sin Pi/4*cos 3Pi/2 Since sin 3Pi/2 = -1; cos 3Pi/2 = 0; cos Pi/4 = V2/2, then, y = (-1)*(V2/2) = -V2/2